# Tripartite measurement uncertainty in Schwarzschild space-time

### v1

@@ Abstract
The effect of Hawking radiation on tripartite measurement uncertainty in a Schwarzschild black hole background is analyzed in this study. Two scenarios are examined: in the first, quantum memory particles approach a Schwarzschild black hole and are positioned near the event horizon, while the particle being measured remains in the asymptotically flat region. In the second scenario, the measured particle moves toward the black hole, and the quantum memories stay in the asymptotically flat region. The study considers two initial quantum states: GHZ and W states. The findings reveal that in both cases, measurement uncertainty increases steadily with rising Hawking temperature. When comparing the GHZ and W states, the GHZ state initially exhibits lower measurement uncertainty at low Hawking temperatures than the W state, indicating greater resilience to Hawking radiation. Additionally, when the quantum memories remain in the asymptotically flat region while the measured particle falls toward the black hole, the uncertainties for GHZ and W states do not align at high temperatures. The GHZ state consistently demonstrates lower measurement uncertainty, showcasing its superior robustness against Hawking radiation.
@@ I Introduction
here, $H(P)=-\sum_{k}p_{k}\log_{2}p_{k}$ represents the Shannon entropy of the measured observable $P\in\{X,Z\}$, where $p_{k}$ is the probability of obtaining measurement outcome $k$. The term $c$ measures the complementarity between the observables and is defined as $c=\max_{\{\mathbb{X},\mathbb{Z}\}}|\langle x_{i}|z_{j}\rangle|^{2}$, with $\mathbb{X}=\{|x_{i}\rangle\}$ and =$\mathbb{Z}=\{|z_{j}\rangle\}$ being the eigenbases of the observables $X$ and $Z$, respectively.
In the last decade, many efforts have been made to generalize and modify this relation [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Uncertainty relations can be understood through a tripartite scenario illustrated by an uncertainty game involving three participants: Alice, Bob, and Charlie. At the start of the game, Alice, Bob, and Charlie share a quantum state $\rho_{ABC}$. In the next step, Alice performs one of two possible measurements, $X$ or $Z$, and then informs Bob and Charlie, who hold the quantum memories $B$ and $C$, respectively, about her measurement choice. If $X$ is measured, Bob’s task is to predict the outcome of this measurement. If $Z$ is measured, then Charlie’s task is to predict the outcome of Alice’s measurement. It has been demonstrated that the tripartite quantum memory entropic uncertainty relation (QM-EUR) can be formulated as [4, 3].
@@
where $c$ is identical to the value defined in Eq. (1). Besides, $S(X|B)$ and $S(Z|C)$ represent the conditional von Neumann entropies of the states
@@
respectively. Physically speaking, $S(X|B)$ quantifies Bob’s uncertainty about the outcome of measurement $X$ given that Bob has access to the quantum memory $B$ and, likewise, $S(Z|C)$ quantifies Charlie’s uncertainty about the outcome of measurement $Z$ given that Charlie has access to the quantum memory $C$.
The study of the influence of the relativistic effect on quantum correlations in curved space-time has been the main topic of many researches [28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Especially, the influence of the Hawking effect on quantum entanglement is studied in the background of a Schwarzschild black hole [36, 37]. It has been shown that in curved space-time, the presence of Hawking radiation can reduce quantum correlations [36, 37, 38, 39]. Also, it is well known that quantum correlations between the quantum memory and the measured particle play an important role in QM-EURs. Note that quantum entanglement between the memory particle and the measured particle can decrease uncertainty. Therefore, it is important to examine how the Hawking effect impacts the tripartite QM-EUR in the context of black holes. While the Hawking effect on QM-EUR has been thoroughly explored in bipartite systems [40, 41, 42, 43], to the best of our knowledge, the tripartite QM-EUR has been investigated in only a few studies [44].
Driven by these observations, we explore the impact of Hawking radiation on the tripartite QM-EUR within Schwarzschild space-time. In our study, we assume that Alice, Bob, and Charlie initially share a generally tripartite entangled state in flat Minkowski space-time. In the next step, we consider two different scenarios: in the first scenario, Bob and Charlie, who hold the quantum memories $B$ and $C$, freely fall towards a Schwarzschild black hole and position themselves near the event horizon, while Alice stays in the asymptotically flat region. In the second scenario, Alice freely falls toward a Schwarzschild black hole and then hovers near the event horizon, while Bob and Charlie remain in the asymptotically flat region. We examine the behavior of the tripartite QM-EUR in these scenarios for two distinct classes of tripartite entangled states: the Greenberger–Horne–Zeilinger (GHZ) state and the W state.
The structure of this Letter is as follows: In Sec. II, we provide a brief overview of the vacuum structure and Hawking radiation for Dirac fields in Schwarzschild space-time. In Sec. III, we analyze the impact of the Hawking effect on the tripartite QM-EUR in Schwarzschild space-time. The final section is dedicated to the conclusion and discussion.
@@ II Dirac fields in Schwarzschild space-time
$\Psi^{I}_{k}=\xi e^{-i\omega u},\quad\Psi^{II}_{k}=\xi e^{+i\omega u},$ (7)
respectively, where $k$ represents the wave vector, $\xi$ is a four-component Dirac spinor, $\omega$ is a monochromatic frequency of the Dirac field. The retarded time $u$ is expressed as $u=t-r_{\ast}$ in which $r_{\ast}=r+2M\ln\frac{r-2M}{2M}$ is the tortoise coordinate.
@@
$|0\rangle_{K}=\alpha|0\rangle^{+}_{I}|0\rangle^{-}_{II}+\beta|1\rangle^{+}_{I}|1\rangle^{-}_{II},$ (8)
@@ III Results
Now, let’s explore the above scenarios for two different initial states.
Figure 3: Measurement uncertainties for both cases when the three qubits are initially prepared in the GHZ state. Upper panels: $U$ versus $\omega$ and $T$. Lower plot: $U$ versus $T$ with $\omega=1$.
@@ III.1 GHZ state
The initial state of the system that has been shared between Alice, Bob, and Charlie is assumed to be a GHZ state
@@
For Case 1, applying (8) and (9), Eq. (10) can be re-expressed in terms of Minkowski modes for Alice and black hole modes for Bob and Charlie as follows:
@@
Using (2) and (12), we obtain the analytical expression of tripartite measurement uncertainty as follows
@@
Using (2) and (15), one can obtain the analytical form of tripartite measurement uncertainty, given by
@@
In Fig. 3, the measurement uncertainties for the GHZ state are calculated based on Eqs. (13) and (16). In the upper panel, this uncertainty is plotted against the monochromatic frequency $\omega$ of the Dirac field and the Hawking temperature $T$ for two distinct scenarios: Case 1 and Case 2. Lower panel of Fig. 3 shows $U$ versus $T$ for $\omega=1$.
@@
If we consider the bottom plot in Fig. 3, one can analyze which of the two cases exhibits a higher $U$ for a fixed $\omega$. In Case 1, Bob and Charlie (quantum memories) are approaching the event horizon, while Alice (the measured particle) remains in the asymptotically flat region. Quantum memories near the event horizon are directly exposed to the intense gravitational effects and Hawking radiation. This exposure is expected to induce greater decoherence and entanglement degradation due to the strong interaction with the thermal radiation emanating from the black hole. Compared with Case 2, where Alice (the measured particle) is near the event horizon, while Bob and Charlie remain in the asymptotically flat region, Case 1 should feature a higher measurement uncertainty $U$ as a function of Hawking temperature $T$, which is demonstrated in Fig. 3.
@@
Figure 4: Measurement uncertainties for both cases when the three qubits are initially prepared in the W state. Upper panels: $U$ versus $\omega$ and $T$. Lower plot: $U$ versus $T$ with $\omega=1$.
@@ III.2 W state
Equipped with (2) and (19), we arrive at
@@
The measurement uncertainties for the W state [see (III.2) and (III.2)] are plotted as a function of the monochromatic frequency $\omega$ of the Dirac field and the Hawking temperature $T$ in Fig. 4. Similar to the GHZ state, for both cases, the measurement uncertainty $U$ decreases with increasing $\omega$ at low temperatures. The decline of $U$ for higher values of $\omega$ at low temperatures again suggests that higher frequency modes are less disruptive to the W state’s entanglement.
@@
In Fig. 4, we also present $U$ as a function of $T$ with a fixed $\omega$. Our observation indicates that the measurement uncertainty behaves differently for the W state compared to the GHZ state at lower temperatures. Specifically, for the W state, Case 1 exhibits less uncertainty at lower temperatures, but as the temperature increases, the uncertainties for both cases tend to converge to the same value.
@@ IV Conclusion
Several studies have investigated quantum correlations in a tripartite system within the context of a Schwarzschild black hole, revealing that their dynamical behaviors are significantly influenced by the Hawking temperature $T$ [36, 37, 38, 39]. These studies demonstrated that the Hawking effect diminishes quantum correlations. Since measurement uncertainty in a QM-EUR is closely linked to the system’s quantum correlations, it is anticipated that it too may be impacted by the Hawking temperature $T$. In this work, the effect of Hawking radiation on the tripartite QM-EUR in Schwarzschild space-time was studied for GHZ and W states. It has been shown that the behaviors of uncertainty depend on the Hawking effect. Specifically, it has been found that the uncertainty increases monotonically with increasing Hawking temperature. As the Hawking temperature $T$ increases, the intensity of Hawking radiation also increases, leading to greater decoherence and increased measurement uncertainty. However, the higher monochromatic frequency $\omega$ of the Dirac field might mitigate some of the uncertainties introduced by the Hawking radiation.

### v2

@@ Abstract
Abstract. The effect of Hawking radiation on tripartite measurement uncertainty in a Schwarzschild black hole background is analyzed in this study. Two scenarios are examined. In the first, quantum memory particles approach a Schwarzschild black hole and are positioned near the event horizon, while the particle being measured remains in the asymptotically flat region. In the second scenario, the measured particle moves toward the black hole, and the quantum memories stay in the asymptotically flat region. This study considers two initial quantum states, namely GHZ and W states. Our findings reveal that in both cases, measurement uncertainty increases steadily with rising Hawking temperature. When comparing the GHZ and W states, the GHZ state initially exhibits lower measurement uncertainty at low Hawking temperatures than the W state, indicating greater resilience to Hawking radiation. Additionally, when the quantum memories remain in the asymptotically flat region while the measured particle falls toward the black hole, the uncertainties for GHZ and W states do not align at high temperatures. The GHZ state consistently demonstrates lower measurement uncertainty, showcasing its superior robustness against Hawking radiation.
@@ I Introduction
where $H(P)=-\sum_{k}p_{k}\log_{2}p_{k}$ represents the Shannon entropy of the measured observable $P\in\{X,Z\}$, in which $p_{k}$ is the probability of obtaining measurement outcome of $k$. The term $c$ measures the complementarity between the observables and is defined as $c=\max_{\{\mathbb{X},\mathbb{Z}\}}|\langle x_{i}|z_{j}\rangle|^{2}$, with $\mathbb{X}=\{|x_{i}\rangle\}$ and =$\mathbb{Z}=\{|z_{j}\rangle\}$ being the eigenbases of the observables $X$ and $Z$, respectively.
In the last decade, many efforts have been made to generalize and modify this relation [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Uncertainty relations can be understood through a tripartite scenario illustrated by an uncertainty game involving three participants: Alice, Bob, and Charlie. At the start of the game, Alice, Bob, and Charlie share a quantum state $\rho_{ABC}$. In the next step, Alice performs one of two possible measurements, $X$ or $Z$, and then informs Bob and Charlie, who hold the quantum memories $B$ and $C$ respectively, about her measurement choice. If $X$ is measured, Bob’s task is to predict the outcome of this measurement. If $Z$ is measured, then Charlie’s task is to predict the outcome of Alice’s measurement. It has been demonstrated that the tripartite quantum memory entropic uncertainty relation (QM-EUR) can be formulated as [4, 3]
@@
where $c$ is identical to the term defined in Eq. (1). Besides, $S(X|B)$ and $S(Z|C)$ represent the conditional von Neumann entropies of the states
@@
respectively, where $\mathbf{I}_{B(C)}$ is the identity operator in Bob (Charlie)’s Hilbert space.
Physically speaking, $S(X|B)$ quantifies Bob’s uncertainty about the outcome of measurement $X$ given that Bob has access to the quantum memory $B$ and, likewise, $S(Z|C)$ quantifies Charlie’s uncertainty about the outcome of measurement $Z$ given that Charlie has access to the quantum memory $C$.
The study of the influence of the relativistic effect on quantum correlations in curved space-time has been the main topic of many researches [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Especially, the influence of Hawking radiation on quantum entanglement is studied in the background of a Schwarzschild black hole [43, 44]. It has been shown that in curved space-time, the presence of Hawking radiation can reduce quantum correlations [43, 44, 45, 46, 47]. Also, it is well known that quantum correlations between the quantum memory and the measured particle play an important role in QM-EURs. Note that quantum entanglement between the memory particle and the measured particle can decrease uncertainty. Therefore, it is important to examine how the Hawking effect impacts the tripartite QM-EUR in the context of black holes. While the Hawking effect on QM-EUR has been thoroughly explored in bipartite systems [48, 49, 50, 51], to the best of our knowledge, the tripartite QM-EUR has been investigated in only a few studies such as Ref. [52].
Driven by these observations, we explore the impact of Hawking radiation on the tripartite QM-EUR within Schwarzschild space-time. In our study, we assume that Alice, Bob, and Charlie initially share a generally tripartite entangled state in flat Minkowski space-time. In the next step, we consider two different scenarios. In the first scenario, Bob and Charlie, who hold the quantum memories $B$ and $C$, freely fall towards a Schwarzschild black hole and position themselves near the event horizon, while Alice stays in the asymptotically flat region. In the second scenario, Alice freely falls toward a Schwarzschild black hole and then hovers near the event horizon, while Bob and Charlie remain in the asymptotically flat region. We examine the behavior of the tripartite QM-EUR in these scenarios for two distinct classes of tripartite entangled states, i.e. the Greenberger–Horne–Zeilinger (GHZ) state and the W state.
The structure of this paper is as follows. In Sec. II, we provide a brief overview of the vacuum structure and Hawking radiation for Dirac fields in Schwarzschild space-time. In Sec. III, we analyze the impact of the Hawking effect on the tripartite QM-EUR in Schwarzschild space-time. The final section is dedicated to the conclusion and discussion.
@@ II Dirac fields in Schwarzschild space-time
$\Psi^{I}_{k}=\xi e^{-i\omega u}\quad\textmd{and}\quad\Psi^{II}_{k}=\xi e^{+i\omega u},$ (7)
respectively, where $k$ represents the wave vector, $\xi$ is a four-component Dirac spinor, and $\omega$ is a monochromatic frequency of the Dirac field. The retarded time $u$ is expressed as $u=t-r_{\ast}$ in which $r_{\ast}=r+2M\ln\frac{r-2M}{2M}$ is the tortoise coordinate.
@@
$|0\rangle_{k}=\alpha|0\rangle^{+}_{I}|0\rangle^{-}_{II}+\beta|1\rangle^{+}_{I}|1\rangle^{-}_{II}$ (8)
and
@@ III Results
Now, let’s explore the above scenarios for two different initial states as mentioned before.
(a)Refer to caption(b)
Refer to caption(c)
Figure 3: Measurement uncertainties for both cases when the three qubits are initially prepared in the GHZ state. Plots (a) and (b): $U$ versus $\omega$ and $T$. Plot (c): $U$ versus $T$ with $\omega=1$.
@@ III.1 GHZ state
The initial state of the system that has been shared between Alice, Bob, and Charlie is assumed to be a GHZ state, given by
@@
For Case 1, by applying (8) and (9), Eq. (10) can be re-expressed in terms of Minkowski modes for Alice and black hole modes for Bob and Charlie as follows:
@@
Using Eqs. (2) and (12), we obtain the analytical expression of tripartite measurement uncertainty as follows
@@
Employing (2) and (15), one can obtain the analytical form of tripartite measurement uncertainty, given by
@@
In Fig. 3, the measurement uncertainties for the GHZ state are calculated based on Eqs. (13) and (16). In Figs. 3(a) and 3(b), this uncertainty is plotted against the monochromatic frequency $\omega$ of the Dirac field and the Hawking temperature $T$ for two distinct scenarios: Case 1 and Case 2. Besides, Fig. 3(c) shows $U$ versus $T$ for $\omega=1$.
@@
If we consider the 2D plot in Fig. 3(c), one can analyze which of the two cases exhibits a higher $U$ for a fixed $\omega$. In Case 1, Bob and Charlie (quantum memories) are approaching the event horizon, while Alice (the measured particle) remains in the asymptotically flat region. Quantum memories near the event horizon are directly exposed to the intense gravitational effects and Hawking radiation. This exposure is expected to induce greater decoherence and entanglement degradation due to the strong interaction with the thermal radiation emanating from the black hole. Compared with Case 2, where Alice (the measured particle) is near the event horizon, while Bob and Charlie remain in the asymptotically flat region, Case 1 should feature a higher measurement uncertainty $U$ as a function of Hawking temperature $T$, which is demonstrated in Fig. 3.
@@
(a)Refer to caption(b)
Refer to caption(c)
Figure 4: Measurement uncertainties for both cases when the three qubits are initially prepared in the W state. Plots (a) and (b): $U$ versus $\omega$ and $T$. Plot (c): $U$ versus $T$ with $\omega=1$.
@@ III.2 W state
Equipped with Eqs. (2) and (19), we arrive at
@@
The measurement uncertainties for the W state [see (III.2) and (III.2)] are plotted as functions of the monochromatic frequency $\omega$ of the Dirac field and the Hawking temperature $T$ in Figs. 4(a) and 4(b). Similar to the GHZ state, for both cases, the measurement uncertainty $U$ decreases with increasing $\omega$ at low temperatures. The decline of $U$ for higher values of $\omega$ at low temperatures again suggests that higher frequency modes are less disruptive to the W state’s entanglement.
@@
In Fig. 4(c), we also present $U$ as a function of $T$ with a fixed $\omega$. Our observation indicates that the measurement uncertainty behaves differently for the W state compared to the GHZ state at lower temperatures. Specifically, for the W state, Case 1 exhibits less uncertainty at lower temperatures, but as the temperature increases, the uncertainties for both cases tend to converge to the same value.
@@ III.3 Comparison between GHZ and W states
The difference between the GHZ and W states can be further understood by considering the geometry of their entanglement. In the GHZ state, all the entanglement is shared globally among all three qubits, meaning the entanglement is strictly three-way [58, 59]. This monogamous nature of the GHZ state ensures that the entanglement is concentrated, creating a strong, unified quantum correlation that is resistant to external disturbances like Hawking radiation. In contrast, the W state’s entanglement is more distributed, allowing two-qubit subsystems to retain some level of entanglement even if one qubit is lost. However, this lack of strong three-way entanglement means the W state does not have the same collective defense against decoherence. As a result, the more dispersed entanglement in the W state leaves it more susceptible to gradual disruptions, such as those caused by low levels of Hawking radiation, leading to higher measurement uncertainty. The GHZ state’s three-way entanglement geometry, being more tightly bound, helps it maintain coherence more effectively across a range of temperatures.
@@ IV Conclusion
Several studies have investigated quantum correlations in a tripartite system within the context of a Schwarzschild black hole, revealing that their dynamical behaviors are significantly influenced by the Hawking temperature $T$ [43, 44, 45, 46, 47]. These studies demonstrated that the Hawking effect diminishes quantum correlations. Since measurement uncertainty in a QM-EUR is closely linked to the system’s quantum correlations, it is anticipated that it too may be impacted by the Hawking temperature $T$. In this work, the effect of Hawking radiation on the tripartite QM-EUR in Schwarzschild space-time was studied for GHZ and W states. It has been shown that the behaviors of uncertainty depend on the Hawking effect. Specifically, it has been found that the uncertainty increases monotonically with increasing Hawking temperature. As the Hawking temperature $T$ increases, the intensity of Hawking radiation also increases, leading to greater decoherence and increased measurement uncertainty. However, the higher monochromatic frequency $\omega$ of the Dirac field might mitigate some of the uncertainties introduced by the Hawking radiation.
@@
Acknowledgments: H.D. acknowledges the support of project VEGA 2/0183/21 (DESCOM).
@@
Code availability: This manuscript has no associated code/software.